OpenStudy (anonymous):

3a=c-a(b-d) solve for a

OpenStudy (anonymous):

3a = c - ab-ad 3a + ab + ad = c a(3+b+d)=c a = c/(3+b+d)

OpenStudy (anonymous):

No medal?

OpenStudy (anonymous):

I don't know where lillianne went but it helped me too lol